ACSman document style

The title pages are easy

Select the ACSman style as the major documentstyle.

\documentstyle{ACSman}

local macros and etc.

\begin{document}

Then enter the title page

\ACSauthor{<your name>}
\ACScopy  {<year>}

\ACStitle {<document title>} \ACSdoc {<document index>} \ACSrev {<revision date>} \ACSabs {<abstract text>}

\maketitle

Then make the contents and etc. with

\tableofcontents
\listoffigures
\listoftables

Begin the text section with

\textpages

and then use \chapter, \section, and etc. in normal LaTEX fashion.

Tools

Be careful of the special symbols

You can make bold text by enclosing it in double quotes "text to be bold" You can make italic text by enclosing it in arrows <text to be italic> While in bold mode you can use \\ to typeset a backslash.

For example, ‘Some "bold" and some <italic> text.’ produces: ‘Some bold

and some italic text.’

Displayed commands are easily formatted with the \<...\> command sequence, which is analogous to \[...\] except that it formats text rather than mathematics. Inside the arrows \\ may be used to seperate lines.